Firm’s total cost function is given as 4Q3 – 40Q2 + 10Q. This firm operates in a perfectly competitive market. What is its profit maximizing level of output?



1
Expert's answer
2021-12-16T11:49:50-0500

"TC=4Q^3-40Q^2+10Q"

"MC=12Q^2-80Q+10"

"MR=0"

Profit maximizing level of output is found when:

"MC= MR"

"12Q^2-80Q+10=0"

Solving the equation above using quadratic formula, we get:

"Q=\frac{80+\sqrt {6400-480}}{24}=6.54"

Profit maximizing level of output is 6.54.